Question 9 (1.5 points) A medical researcher wants to construct a 99% confidence interval for the proportion of knee replacement surgeries that result in complications. An article in the Journal of Bone and Joint Surgery suggested that approximately 8% of such operations result in complications. Using this estimate, what sample size is needed so that the confidence interval will have a margin of error of o.023 2 Write only an integer as your answer. Your Answer Answet Save Question 10 (1.5 points) A researcher wants to construct a 99% confidence interval for the proportion of people who changed jobs in the past year What sample size is needed so that the confidence interval will have a margin of error of 0.031 7 Write only an integer as your answer. Your Answer; Answer Save Question 11 (1.5 points) A pollster wants to construct a 95% confidence interval for the proportion of adults who believe that economic conditions are getting better, A Gallup poll taken in July 2010 estimates this proportion to be o.33. Using this estimate, what sample size is needed s?that the confidence interval will have?margin of error of 0.026 ? Write only an integer as your ansver Your Answers
Answer #1

9 ) Determination of sample size necessary for estimating population proportion :

n=\left ( \frac{Z}{E} \right )^2\times p(1-p)

where , Z= 2.576\ , p=0.08\ , E=0.023

Required sample size :

n=923.24\approx 924

10 ) Determination of sample size necessary for estimating population proportion :

n=\left ( \frac{Z}{E} \right )^2\times p(1-p)

where , Z= 2.576\ , p=0.50\ , E=0.031

Required sample size :

n=1726.27\approx 1727

11 ) Determination of sample size necessary for estimating population proportion :

n=\left ( \frac{Z}{E} \right )^2\times p(1-p)

where , Z= 1.96\ , p=0.33\ , E=0.026

Required sample size :

n=1256.48\approx 1257